OpenStudy (anonymous):

What is the value of x? Round to the nearest tenth. A. 18.4 B. 19.2 C. 20.4 D. 25.1

OpenStudy (alexandervonhumboldt2):

equation?

OpenStudy (studygurl14):

What's the equation? @beklekchek

OpenStudy (alexandervonhumboldt2):

it akss to sing in

OpenStudy (anonymous):

Welcome to openstudy!

OpenStudy (alexandervonhumboldt2):

just type your equation here

OpenStudy (anonymous):

whoops thats the wrong page

OpenStudy (alexandervonhumboldt2):

asks to sing in.........

OpenStudy (anonymous):

um there is not an equation the picture is of a right triangle with sides 12 and 22 and you need to find the value of x

OpenStudy (alexandervonhumboldt2):

so

OpenStudy (alexandervonhumboldt2):



OpenStudy (alexandervonhumboldt2):

like this?

OpenStudy (alexandervonhumboldt2):

ifyes then use pyth theore,: 22^2 + 12^2 = x^2

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

really that simple?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

Then you need to find the square root of the answer

OpenStudy (alexandervonhumboldt2):

just use cacculator if you cant count 22^2 + 12^2 = side^2 dont forget to take sqauare root!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

did i medal you @alexandervonhumboldt2?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (anonymous):

how do i?

OpenStudy (alexandervonhumboldt2):

you cn medal by clicking best responce abd fan by clicking alexandervonhumboldt2 and then fan

OpenStudy (anonymous):

did it! is the answer 25.1

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

even though yall arent replying thanks for responding so quickly

OpenStudy (anonymous):

@AlexandervonHumboldt2 @studygurl14